F_{elastica} = k.e (1)

F_{elastica} k e e = (e_{f}−

e_{i}) F_{elastica}

∑ \overrightarrow{M_{O}} = \overrightarrow{0} (2)

?? 45^{∘} ?? θ_{2}

pass_{t}rc2_{1}2[P osiesdocentromassa : inicialefinal]P osiesdocentromassa : inicialefinal.

?? r_{cm}

\overrightarrow{F}_{elstica} r_{cm} r_{cm} 109.99mm 112.00mm soft- ware ?? pass_{d}cl11[Diagramascorpolivre]Diagramascorpolivre.P osioinicialefinaldotronco.

??

−r_{elstico}×F_{elstico}+r_{cm}×F = 0

−r_{elstico}.F_{elstico}. sin(θ)+r_{cm}(γ). cos(α).F = 0

F_{elstico} = \frac{r_{cm}(γ). cos(α).F}{r_{elstico}. sin(θ)}

r_{cm}(β) β

0^{∘} 45^{∘} α 89.53^{∘} 41.56^{∘} θ 67.60^{∘} 38.55^{∘} ??

k = \frac{r_{cm}(γ).F. cos(α)}{r_{elstico}. sin(θ).e(γ)}

e = (e_{f}−

e_{i})

37.76 73.40 e_{i} 40mm k 16.25 934.70 N.m^{−1} k

\overrightarrow{F}_{elstica} ≈ 0.47^{∘} ?? 45^{∘} ?? θ_{1}

pass_{t}rc1_{1}2[P osiesdocentromassa : inicialefinal]P osiesdocentromassa : inicialefinal.

\overrightarrow{F}_{elstica} ?? pass_{d}cl21[P osiesdocentromassa :

inicialefinal]P osiesdocentromassa : inicialefinal. ??

−r_{elstico}×F_{elstico}−r_{cm}×F = 0

−r_{elstico}.F_{elstico}. sin(β)−r_{cm}(γ). cos(α).F = 0

F_{elstico} = −\frac{r_{cm}(γ). cos(α).F}{r_{elstico}. sin(β)}

r_{cm}(γ) γ

0^{∘} 45^{∘}

F_{elstico} = F_{e}_{1}+F_{e}_{2}. cos(θ)

??

F_{elstico} = k_{e}_{1}.e_{1}(γ)+k_{e}_{2}.e_{2}(γ). cos(θ)

k_{e}_{1} =

k_{e}_{2}

F_{elstico} = k.(e_{1}(γ)+e_{2}(γ). cos(θ))

e_{1} = (e_{1}_{f}−

e_{1}_{i}) e_{2} =

(e_{2}_{f}− e_{2}_{i})